Need with help understanding gauss elimination in a simple way.

−3x[2] + 7x[3] = 4
x[1] + 2x[2] − x[3] = 0
5x[1] − 2x[2] = 3

Use Gauss elimination with partial pivoting to solve for the x’s. As part of the computation, Calculate the determinant.
Answer #1

ox, -322 -3n2+ 743 = 4 *, +2^2 - 43 = 0 5x, -202=3 In partial pivoting the last equation is taken . are equation 0 as coefficputting in in equation ④ 5n, -2n2 =3 :: 5x, = 3 + 2x 2009 5% = 3 + you 5*, = 247 69 n = 247 345 ② in writing the equation o,